# Adoption of Watermarking Measures for AI-Generated content and Implications under the EU AI Act

### v2

@@ Abstract.
AI-generated images have become so good in recent years that individuals often cannot distinguish them any more from “real” images. This development, combined with the rapid spread of AI-generated content online, creates a series of societal risks, particularly with the emergence of “deep fakes” that impersonate real individuals. Watermarking, a technique that involves embedding information within images and other content to indicate their AI-generated nature, has emerged as a primary mechanism to address the risks posed by AI-generated content. Indeed, watermarking and AI labelling measures are now becoming a legal requirement in many jurisdictions, including under the 2024 European Union AI Act. Despite the widespread use of AI image generation systems, the current status of the implementation of such measures remains largely unexamined. Moreover, the practical implications of the AI Act’s watermarking and labelling requirements have not previously been studied.
The present paper therefore both provides an empirical analysis of 50 widely used AI systems for image generation, embedded into a legal analysis of the AI Act. In our legal analysis, we identify four categories of generative AI image deployment scenarios relevant under the AI Act and outline how the legal obligations apply in each category. In our empirical analysis, we find that only a minority number of AI image generators currently implement adequate watermarking (38%) and deep fake labelling (8%) practices. In response, we suggest a range of avenues of how the implementation of these legally mandated techniques can be improved, and publicly share our tooling for the easy detection of watermarks in images.
@@ 1. Introduction
To align financial and societal incentives better, companies face increasing legal requirements in the AI space. The 2024 EU AI Act mandates two key measures to mitigate the risks posed by AI-generated content: (1) the embedding of machine-readable markings in all AI-generated outputs to facilitate automated detection of synthetic content and (2) visibly disclosing the artificial origin of AI-generated “deep fakes” (European Union, 2024). In case of non-compliance, companies face high potential fines: up to 15 million Euros or 3% of a company’s global annual turnover. These rules will start to apply from 1 August 2026, and their implementation could prove to be a foundational step towards ensuring trust and mitigating the risks of generative AI content. However, significant ambiguities exist around the practical application of these requirements. This includes questions on how to allocate responsibilities along the complex generative AI supply chain, how to assess the adequacy of existing technical solutions, and how to interpret definitions used by the AI Act (e.g., the definition of a ‘deep fake’) (Dewitte, 2024; Engler and Renda, 2022; Meding and Sorge, 2024).
@@
  • Legal analysis and clarification of the EU AI Act: We clarify the applicability of the AI Act’s transparency requirements in different common deployment scenarios of generative AI text-to-image systems. These insights contribute to a successful implementation of the EU AI Act in practice and to the design of AI content transparency regulation beyond the EU.
@@
Since the AI Act’s transparency requirements are only applicable from August 2026 onwards, the purpose of our empirical study is not to evaluate compliance. Rather, we aim to present and discuss the current status quo of the implementation of machine-readable (water)marking and disclosure solutions, and the challenges associated with implementing and enforcing legal requirements (such as in the AI Act) related to such solutions.
Structure. The remainder of this paper is structured as follows. Section 2 reviews relevant literature on existing technical solutions for marking and labelling AI-generated image content and introduces the general framework of the EU AI Act. Section 3 presents a legal analysis of the AI Act’s transparency provisions, focusing on how these requirements apply in common deployment scenarios of generative AI systems. Section 4 details the methodology used in the empirical analysis. Section 5 presents the results of this analysis. Finally, we discuss the implications of these findings for the (implementation of the) AI Act— and beyond.
@@ 2. Background
2.1. Techniques for marking and labelling AI-generated image content
@@
Several developers are starting to use invisible machine-based watermarking solutions in their image generation models. Meta claims to use an invisible watermark (next to visible marking and metadata embeddings) that is applied using a deep learning model, but they do not share the corresponding model that can be used to detect this watermark (Meta AI, 2023). Meta has also worked with other researchers on watermarking solutions that can be embedded during the image generation process, which could prove valuable for open-source models (Fernandez et al., 2023; Meta AI, 2024). Google adds an invisible watermark into the pixels of AI-generated images from their Imagen models (Adobe, 2025), while also offering a related open-source detection tool (Google Cloud, nd). The open-source model developers Stability AI and Black Forest Labs (together with their partners) also integrate invisible watermarking techniques in (some of) their models (Stability AI, ndb; AI, 2024). They use an open-source watermarking library that includes both a watermarking encoding and decoder algorithm (Stability AI, nda; ShieldMnt, nd; Black Forest Labs, 2025). Additionally, several organisations are offering API-based solutions or open-source tools that allow others to embed post-generation invisible watermarks to their images (IMATAG, nd; Hugging Face, ndb).
@@
2.2. Approach and scope of the EU AI Act
The EU AI Act was adopted in 2024 to foster safe and responsible AI development and deployment in the EU. Although other jurisdictions, such as South Korea (South Korean National Assembly, 2024), have also introduced AI laws that contain relevant transparency rules for generative AI content, the EU AI Act is the most comprehensive and wide-ranging regulatory framework for AI to date and is therefore the reference for this paper.
@@ 2.2. Approach and scope of the EU AI Act
AI Act distinguishes between providers (those developing and/or marketing AI systems or GPAI models under their own name/trademark) and deployers (those deploying AI systems to end-users under their own authority). Most obligations fall onto providers, with deployers facing some limited requirements in the case of deploying high-risk systems or certain AI systems with transparency risks. End-users (individuals or organisations that interact with — or are affected by — an AI system) do not face any legal obligations. Most of the AI Act’s provisions, including the transparency requirements, take effect on August 1, 2026, with fines up to €15 million, or 3% of annual global turnover in case of non-compliance with the transparency requirements.
@@ 3. Legal analysis: Watermarking obligations under the EU AI Act
Section 3.1 outlines the relevant provisions for marking and labelling of AI-generated content as set out in Article 50 of the AI Act. While these requirements may appear straightforward, their practical application can prove difficult due to the complex supply chains behind generative AI systems. To address these complexities, Section 3.2 introduces four common deployment scenarios of generative AI systems and analyses how the requirements apply in each scenario. These scenarios subsequently serve as our basis for the empirical study of Section 4.
@@ 3.1. The AI Act’s rules for watermarking and labelling AI-generated content
Further to this technical marking requirement, Article 50(4) mandates that deployers of generative AI systems that “generate or manipulate image, audio or video content constituting a deep fake have to disclose that the content has been artificially generated or manipulated”. Article 3(60) thereby defines a deep fake as an “AI-generated or manipulated image, audio or video content that resembles existing persons, objects, places, entities or events and would falsely appear to a person to be authentic or truthful”. The different methods that can be used for implementing this disclosure are not specified. Article 50(5) however does state that the information referred to in both these transparency rules “shall be provided to the natural persons concerned in a clear and distinguishable manner at the latest at the time of the first interaction or exposure”. For the purpose of this study, we interpret this as a requirement to include a visible watermark, label or disclosure message directly in the output of a generative AI system, instead of adding a separate disclosure message that is not included in the image (e.g., a text label in the interface of the system). This ensures that the marking remains visible when the image is shared without any manipulations.
The transparency rules of Article 50 apply only to AI systems (e.g., apps or web tools), meaning that the developers and providers of the AI models behind generative AI applications are not required to implement any of the transparency requirements. As also written in the AI Act, “AI models, including General-Purpose AI models, require the addition of further components, such as for example a user interface, to become AI systems” (Recital 97). This is commonly achieved through a web interface, a mobile application, or a desktop client.
The requirements for the separate category of GPAI-models do not include any measures directly aimed at AI-generated content. The third draft of the Code of Practice for GPAI models does, however, currently include a commitment for signatories of the Code to use methods such as watermarks and digital fingerprinting for post-market monitoring of systemic risks and the tracing of incidents related to the use of their model ((European Commission, 2024)).
Important to note is that in contrast to some other AI Act requirements, providers of AI systems released under free and open-source licenses are not exempt from the transparency requirements of Article 50 (Article 2(12)). Moreover, the AI Act transparency provisions are likely to apply to a large number of non-EU providers too, as most publicly available generative AI apps and websites can and are being used within the EU.
@@ 3.2. How the rules apply in practice
Generative AI applications, including image generators, are the product of a complex and diverse supply chain, involving different interconnected steps and a large variety of different participants (Dewitte, 2024; Lee et al., 2023). A simplified version of the generative AI image supply chain can be defined as consisting of four primary players performing different actions: 1) base model developers, 2) downstream developers, 3) system providers, and 4) system deployers. Figure 1 illustrates how the AI Act’s rules would apply along this simplified version of the generative AI supply chain.
@@
In practice, however, the distinction between model developers, system providers, and deployers is less clear-cut and the implementation of the transparency requirements is therefore not strictly divided between these roles. To clarify the application of the AI Act’s transparency rules in practice, we identify four common deployment scenarios of generative AI systems and analyse how the rules apply in each case. These four scenarios thereby serve as the foundation for the measurement study of Section 4.
@@
These organisations thus become both the provider and deployer of the generative AI system under the AI Act, which will require compliance with both transparency requirements. The dual role of model and system developer would allow such organisations to implement robust watermarking techniques in the model development phase (such as embedding initial noise patterns in diffusion-based models).
@@
Some organisations in this end-to-end category also operate large (social) media or image platforms (e.g., Google, X), which requires these providers to comply with the systemic risk mitigation rules under the EU’s Digital Services Act. Furthermore, certain GPAI models integrated in these end-to-end systems may fall under the AI Act’s category of GPAI models with systemic risk, which would require the implementation of additional systemic risk mitigation measures.
@@
Many generative AI image systems are built by leveraging (easy-to-use) APIs from large-scale GPAI model providers. Developers then integrate the model APIs into user-friendly interfaces, such as mobile apps or web tools, hereby creating the generative AI system as defined under the AI Act. This approach allows organisations to use high-quality models without needing the expertise or infrastructure to build and run them. Providers of such systems often also offer additional functionalities to end-users, like the option to choose between different models or the addition of system prompts to help users improve their creations.
Organisations leveraging models via APIs are typically both the provider and deployer of the generative AI system and, thus, will need to comply with both transparency requirements. Compliance can be achieved by relying on features (such as invisible watermarks) that are already built into the model or offered by model providers as an API-parameter option. Otherwise, organisations would have to implement their own measures after the images are created via the API. For example, through the addition of post-processing watermarks, or the embedding of metadata. Applying visible disclosure measures solely to deep fake images could thereby provide a significant challenge (especially for smaller-scale organisations), as this would require a separate (NLP-based) solution capable of specifically detecting deep fake prompts. A simpler approach could involve applying visible disclosures to all generated images, although this may negatively impact the user experience and business model of a provider.
In some cases, providers of model APIs can also offer pre-built code or templates for creating a user interface alongside the API. These model providers could then potentially be considered as system providers under the AI Act, thus bearing responsibility for the machine-readable marking requirement. However, as guidance is still to be created on the distinction between providers and deployers and such situations are relatively rare in practice, we do not analyse this case in this study.
@@
By providing a platform and tools for storing, testing, sharing and deploying AI models, Hugging Face has an important role in the generative AI ecosystem. Many AI image model developers share their open-source models on Hugging Face and thereby allow others to download, use and potentially fine-tune these models for specific applications. Hugging Face thereby also offers so-called interactive API inference widgets, which are pre-configured interfaces added to a model page on Hugging Face that directly allow end-users to use a model to generate images, as shown in Figure 2 (Hugging Face, nda). Hugging Face provides the infrastructure for this tool, and by default, the widgets are enabled by Hugging Face for publicly available models. Model owners do retain full control of the model and can adjust the settings of these widgets.
Figure 2. Hugging Face’s Inference API widget (bottom right), as found on the popular stable-diffusion model page
@@
We include this category because in terms of the AI Act, it is not trivial to decide who would bear the responsibility for complying with the AI Act’s requirements in this case. It could be argued that Hugging Face, which offers the user interface but has no control over the model, would be the provider and the deployer in some of these cases. Plus, we include this category to also study whether content of popular open-source image models, like the models from Stability AI, also includes pre-baked watermarking solutions when generated via such Hugging Face tools.
@@
The final category includes organisations that use (finetuned) AI models of other organisations in their own system but deploy them under their own name and trademark, without disclosing the underlying models used. Organisations in this category would then classify as both the provider and deployer of a generative AI system, requiring compliance with both transparency obligations. We include this category as there are many websites and mobile apps, often listed in non-EU countries, that offer generative AI systems that fall under this category.
Providers of such systems could hereby rely on downloaded versions of open-source models. This allows providers to finetune these models and possibly also share them further via APIs or open-source repository platforms, but also requires infrastructure to host and run these models. Or they can use API model access without disclosing or crediting the provider of the API. While API license terms typically require crediting the model developer, exceptions exist, or organisations may disregard the license terms.
The next section explores whether and how watermarking solutions are currently being implemented in practice, using the four deployment scenarios outlined above as the basis for this analysis. In practice, however, distinguishing between these categories can prove challenging, due to the many ways in which generative AI models and systems are developed and shared.

### v3

# Missing the Mark: Adoption of Watermarking for Generative AI Systems in Practice and Implications under the new EU AI Act

@@ Abstract.
AI-generated images have become so good in recent years that individuals often cannot distinguish them any more from “real” images. This development, combined with the rapid spread of AI-generated content online, creates a series of societal risks. Watermarking, a technique that involves embedding information within images and other content to indicate their AI-generated nature, has emerged as a primary mechanism to address the risks posed by AI-generated content. Indeed, watermarking and AI labelling measures are now becoming a legal requirement in many jurisdictions, including under the 2024 European Union AI Act. Despite the widespread use of AI image generation systems, the practical implications and the current status of implementation of these measures remain largely unexamined.
The present paper therefore provides both an empirical and a legal analysis of these measures. In our legal analysis, we identify four categories of generative AI deployment scenarios and outline how the legal obligations could apply in each category. In our empirical analysis, we find that only a minority number of AI image generators currently implement adequate watermarking (38%) and deep fake labelling (18%) practices. In response, we suggest a range of avenues of how the implementation of these legally mandated techniques can be improved, and publicly share our tooling for the detection of watermarks in images. [^1] This paper was accepted to the 2025 ACM CS & Law Work-in-Progress track.
@@ 1. Introduction
To align financial and societal incentives better, companies face increasing legal requirements in the AI space. The 2024 EU AI Act mandates two key measures to mitigate the risks posed by AI-generated content: (1) the embedding of machine-readable markings in all AI-generated outputs to facilitate automated detection of synthetic content and (2) visibly disclosing the artificial origin of AI-generated “deep fakes” (European Union, 2024). In case of non-compliance, organisations face high potential fines: up to 15 million Euros or 3% of a company’s global annual turnover. These rules will start to apply from 1 August 2026, and their implementation could prove to be a foundational step towards ensuring trust and mitigating the risks of generative AI content. However, significant ambiguities exist around the practical application of these requirements. This includes questions on how to allocate responsibilities along the complex generative AI supply chain, how to assess the adequacy of existing technical solutions, and how to interpret definitions used by the AI Act (e.g., the definition of a ‘deep fake’) (Dewitte, 2024; Engler and Renda, 2022; Meding and Sorge, 2024).
@@
  • Legal analysis and clarification of the EU AI Act: We analyse the applicability of the AI Act’s transparency requirements in different common deployment scenarios of generative AI text-to-image systems. These insights can contribute to a successful implementation of the EU AI Act in practice and inform the design of AI content transparency measures beyond the EU.
@@
Since the AI Act’s transparency requirements are only applicable from August 2026 onwards, the purpose of our empirical study is not to evaluate compliance. Rather, we aim to present and discuss the current status quo of the implementation of machine-readable (water)marking and visible disclosure solutions for AI-generated image content, and the challenges associated with implementing and enforcing legal requirements (such as in the AI Act) related to such solutions.
Structure. The remainder of this paper is structured as follows. Section 2 reviews relevant literature on existing technical solutions for marking and labelling AI-generated image content and introduces the general framework of the EU AI Act. Section 3 presents a legal analysis of the AI Act’s transparency provisions, focusing on how these requirements could apply in different common deployment scenarios of generative AI systems. Section 4 details the methodology used in the empirical analysis. Section 5 presents the results of this analysis. Finally, we discuss the implications of these findings for the (implementation of the) AI Act— and beyond.
@@ 2. Background
2.1. Technical background: Techniques for marking and labelling AI-generated image content
@@
Several developers are starting to use invisible machine-based watermarking solutions in their image generation models. Meta claims to use an invisible watermark (next to visible marking and metadata embeddings) that is applied using a deep learning model, but they do not share the corresponding model that can be used to detect this watermark (Meta AI, 2023). Meta has also worked with other researchers on watermarking solutions that can be embedded during the image generation process, which could prove valuable for open-source models (Fernandez et al., 2023; Meta AI, 2024). Google adds an invisible watermark into the pixels of AI-generated images from their Imagen models (Adobe, 2025), while also offering a related open-source detection tool (Google Cloud, nd). The open-source model developers Stability AI and Black Forest Labs (together with their partners) also integrate invisible watermarking techniques in (some of) their models (Stability AI, ndb; AI, 2024). They use an open-source watermarking library that includes both a watermarking encoding and decoder algorithm (Stability AI, nda; ShieldMnt, nd; Black Forest Labs, 2025). Additionally, several organisations are offering API-based solutions or open-source tools that allow others to embed post-generation invisible watermarks to their images (IMATAG, nd; Hugging Face, ndc).
@@
2.2. Legal background: Approach and scope of the EU AI Act
The EU AI Act was adopted in 2024 to foster safe and responsible AI development and deployment in the EU. Although other jurisdictions, such as South Korea (South Korean National Assembly, 2024) and China (Cyberspace Administration of China, 2025), have also introduced AI laws that contain relevant transparency rules for generative AI content, the EU AI Act is the most comprehensive and wide-ranging regulatory framework for AI to date and is therefore the reference for this paper.
@@ 2.2. Legal background: Approach and scope of the EU AI Act
The AI Act distinguishes between providers (those developing and/or marketing AI systems or GPAI models under their own name/trademark) and deployers (those deploying AI systems to end-users under their own authority). Most obligations fall onto providers, with deployers facing some limited requirements in the case of deploying high-risk systems or certain AI systems with transparency risks. End-users (individuals or organisations that interact with — or are affected by — an AI system) do not face any legal obligations. Most of the AI Act’s provisions, including the transparency requirements, take effect on August 1, 2026, with fines up to €15 million, or 3% of annual global turnover in case of non-compliance with the transparency requirements. The AI Act distinguishes between providers (those developing and/or marketing AI systems or GPAI models under their own name/trademark) and deployers (natural or legal persons using an AI system under their own authority, except where this use is for a personal non-professional activity. Most obligations fall onto providers, with deployers facing some limited requirements in the case of deploying high-risk systems or certain AI systems with transparency risks. Most of the AI Act’s provisions, including the transparency requirements, take effect on August 1, 2026, with fines up to €15 million, or 3% of annual global turnover in case of non-compliance with the transparency requirements.
@@ 3. Legal analysis: Watermarking obligations under the EU AI Act
Section 3.1 outlines the relevant provisions for marking and labelling of AI-generated content as set out in Article 50 of the AI Act. While these requirements may appear straightforward, their practical application can prove difficult due to the complex supply chains behind generative AI systems. To address these complexities, Section 3.2 introduces four common deployment scenarios of generative AI systems and analyses how the requirements could apply in each scenario. These scenarios subsequently serve as our basis for the empirical study of Section 4.
@@ 3.1. The AI Act’s rules for watermarking and labelling AI-generated content
Further to this technical marking requirement, Article 50(4) mandates that deployers of generative AI systems that “generate or manipulate image, audio or video content constituting a deep fake have to disclose that the content has been artificially generated or manipulated”. Article 3(60) thereby defines a deep fake as an “AI-generated or manipulated image, audio or video content that resembles existing persons, objects, places, entities or events and would falsely appear to a person to be authentic or truthful”. However, what constitutes as a deep fake under this definition is still open to interpretation and far from a trivial task (Meding and Sorge, 2024).
Moreover, the different methods that can be used for implementing this disclosure are not specified. Article 50(5) however does state that the information referred to in both these transparency rules “shall be provided to the natural persons concerned in a clear and distinguishable manner at the latest at the time of the first interaction or exposure”. For this study, we interpret this as a requirement to include a visible watermark, label or disclosure message directly in the output of a generative AI system, instead of adding a separate disclosure message that is not included in the image (e.g., a text label in the interface of the system). This ensures that the marking remains visible when the image is shared without any manipulations.
The transparency rules of Article 50 apply only to AI systems (e.g., apps or web tools), meaning that the developers and providers of the AI models behind generative AI applications are not required to implement any of these requirements. As also written in the AI Act, “AI models, including General-Purpose AI models, require the addition of further components, such as for example a user interface, to become AI systems” (Recital 97). This is commonly achieved through a web interface, a mobile application, or a desktop client.
The requirements for the separate category of GPAI-models do not include any measures directly aimed at AI-generated content. The adopted Code of Practice for GPAI models does, however, currently include a commitment for signatories of the Code to use methods such as watermarks and digital fingerprinting for post-market monitoring of systemic risks and the tracing of incidents related to the use of their model (European Commission, 2025a).
Important to note is that in contrast to some other AI Act requirements, providers of AI systems released under free and open-source licenses are not exempt from the transparency requirements of Article 50 (Article 2(12)). Moreover, the AI Act transparency provisions are likely to apply to a large number of non-EU providers too, as most publicly available generative AI apps and websites can and are being used within the EU (Article 2(1)).
@@ 3.2. How the rules apply in practice
Generative AI applications, including image generators, are the product of a complex and diverse supply chain, iinvolving a wide range of different actors (Dewitte, 2024; Lee et al., 2023). A simplified version of the supply chain for generative AI image supply chain can be described as consisting of four primary players, each performing different functions: 1) base model developers, 2) downstream model developers, 3) system (interface) providers, and 4) system deployers. Figure 1 illustrates how the AI Act’s rules would apply along this simplified supply chain.
@@
In practice, however, the distinction between model developers (or model providers), system providers, and and deployers is less clear-cut. Consequently, and the implementation of the AI Act’s transparency requirements measures is are therefore not strictly divided between these roles and remains subject to interpretation. The European Commission is expected to publish guidelines to further clarify the definitions and scope of the obligations in article 50 (European Commission, 2025b).
One example of this ambiguity is the use of the term deployer in the deep fake disclosure requirement (Barberá, 2024). The AI Act defines deployers as those using an AI system under their own authority (except when this use is in the case of personal non-professional activities). Under this definition, deployers could be interpreted as the ‘end-users’ of generative AI systems, who query a system to generate content that is used in a professional activity (e.g., organisations generating AI content for a marketing campaign), which then requires a deep fake disclosure. This interpretation would align with terminology used in related EU product regulations, such as the Machinery Directive, which uses the term professional user to refer to a natural person operating a machinery product in a professional context (European Parliament and Council, 2023), or the Medical Device Regulations, where user refers to health care professionals or laypeople using the device (European Parliament and Council, 2017a, b).
However, given that the AI Act does not use or define a similar a term, and considering the nature of AI systems, and how the term deployer is used in the high-risk provisions, deployer could also be interpreted as to refer to the entity that deploys a generative AI system under its authority to end-users (e.g., via an app or website). In that case, such an entity also uses an existing system under its authority, although the nature of this usage is obviously different from that of end-users. Given the potential risks associated with generative AI systems, and the likely enforcement challenges for the disclosure requirement, it would arguably be more practical to implement disclosure measures at the system level (such as automatically applied labels within an image, or by giving end-users the option to add a label), rather than expecting professional end-users to apply such labels manually after image generation. Accordingly, and for the purposes of this study, we adopt the latter interpretation when referring to the deployer role.
To further illustrate the challenges of applying the AI Act’s transparency requirements in practice, we identify four common deployment scenarios of generative AI systems and analyse how the relevant rules may apply in each case. These four scenarios thereby serve as the foundation for the measurement study in Sections 4 and 5.
@@
These organisations thus become both the provider and deployer of the generative AI system under our interpretation of the AI Act, which requires compliance with both transparency requirements. The dual role of model and system developer would allow such organisations to implement robust watermarking techniques in the model development phase (such as embedding initial noise patterns in diffusion-based models).
@@
Some organisations in this end-to-end category also operate large (social) media or image platforms (e.g., Google, X), which requires these providers to comply with the systemic risk mitigation rules under the EU’s Digital Services Act, including measures linked to generative AI (European Commission, 2024). Furthermore, certain GPAI models integrated in these end-to-end systems may fall under the AI Act’s category of GPAI models with systemic risk, which would require the implementation of additional systemic risk mitigation measures.
@@
Many generative AI image systems are built by leveraging (easy-to-use) APIs from large-scale GPAI model providers and integrating these APIs into user-friendly interfaces, such as mobile apps or web tools, hereby creating the generative AI system as defined under the AI Act. This approach allows organisations to use high-quality models without needing the expertise or infrastructure to build and run them. Developers of such systems often also offer additional functionalities to end-users, like the option to choose between different models or the addition of system prompts to help users improve their creations.
Organisations leveraging models via APIs would then likely be considered both the provider and deployer of the generative AI system and, thus, will need to comply with both transparency requirements. Compliance can be achieved by relying on features (such as invisible watermarks) that are already built into the model or offered by model providers as an API-parameter option. Otherwise, organisations would have to implement their own measures after the images are created via the API. For example, through the addition of post-processing watermarks, or the embedding of metadata. Applying visible disclosure measures solely to deep fake images could thereby provide a challenge (especially for smaller-scale organisations), as this would require a separate (NLP-based) solution capable of specifically detecting deep fake prompts. A simpler approach could involve applying visible disclosures to all generated images, although this may negatively impact the user experience and business model of a provider.
In some cases, providers of model APIs can also offer pre-built code or templates for creating a user interface alongside the API. These model providers could then potentially be considered as system providers under the AI Act, thus bearing responsibility for the machine-readable marking requirement. However, as guidance is still to be created on the distinction between providers and deployers, we do not further analyse this case in this study.
@@
By providing a platform and tools for storing, testing, sharing and deploying AI models, the Hugging Face has an important role in the generative AI ecosystem. Many AI image model developers share their open-source models on the Hugging Face platform and thereby allow others to download, use and potentially fine-tune these models for specific applications.
Hugging Face thereby also offers so-called interactive API inference widgets, which are pre-configured interfaces added to a model page on Hugging Face that directly allow end-users to use a model to generate images, as shown in Figure 2 (Hugging Face, nda, ndb). Hugging Face works with inference providers who provide the infrastructure for this tool, and model owners can adjust the settings of these widgets.
Figure 2. Example of Hugging Face’s Inference widget (bottom right), as found on the popular stable-diffusion XL-1.0 model page. This widget allows end-users to directly generate images from a model; it is yet unclear who in this case bears responsibility for implementing the AI Act’s transparency requirements.
@@
We include this category because in terms of the AI Act, it is not trivial to decide who would bear the responsibility for complying with the AI Act’s requirements in these cases. It could be argued that Hugging Face or its inference providers, who offer the platform and infrastructure, but have little control over the models or API’s, would be the provider and/or deployer in some of these cases. We also include this category to study whether content of popular open-source image models, like the models from Stability AI, also includes pre-baked watermarking solutions when generated via such Hugging Face tools.
@@
The final category includes organisations that use (finetuned) AI models of other organisations in their own system and deploy them under their own name and trademark (via apps or websites), without disclosing the underlying models used. Organisations in this category would then again likely classify as both the provider and deployer of a generative AI system, requiring compliance with both transparency obligations. We include this category as there are many websites and mobile apps, often listed in non-EU countries, that offer generative AI systems that fall under this category.
Providers of such systems could thereby rely on downloaded versions of open-source models, which allows for finetuning and possibly also sharing or selling the models further (via APIs or open-source repository platforms), which does require infrastructure to host and run these models. Or they can use API model access without disclosing or crediting the provider of the API. While API license terms typically require crediting the model developer, exceptions exist, or organisations may disregard the license terms.
The next section explores whether and how machine-readable marking and deep fake labelling solutions are currently being implemented in practice, using the four deployment scenarios outlined above as the basis for this analysis. In practice, however, distinguishing between these categories can prove challenging, due to the many ways in which generative AI models and systems are developed and shared.